OpenStudy (anonymous):

Best Rentals charges a daily fee for renting its cars. Barney was charged $105 for 3 days and 300 miles, while Mary was charged $195 for 5 days and 600 miles. What does Best Rental charge per day?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

say x is charge/day y is charge/300mile 3x+y=105 5x+2y=195

OpenStudy (anonymous):

still working on it!!

OpenStudy (anonymous):

your answer is good I am just doing the equation

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

medal would help too!

OpenStudy (anonymous):

Shadow could you break it down for me, I still don't get it

OpenStudy (anonymous):

say x is charge/day --------------rent for a car y is charge/300mile ---------charge for mileage Barney rented for 3 days and drove 300 miles 3x+y=105 Mary rented for 5 days and drove 600 miles 5x+2y=195

OpenStudy (anonymous):

ok so far I got that part

OpenStudy (anonymous):

then multiply first equation by 2 and subtract second equation 6x + 2y = 210 -5x -2y = -195 y terms eliminate x= 15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

let d be the cost per day to rent the car and m be the cost per mile driven. Solve the following simultaneous equations for d and m: 3d+300m=105,5d+600m=192 d =$15/day, m=$1/5 per mile or 20 cents per mile